The moment of interia of a triangular lamina of mass M and height H about its base

Grade:11

2 Answers

harshin nechooli
5 years ago
Main asking for moment of inertia you should first specify the axis of rotation I think it is on the base.        the     MHsq ÷6
Khimraj
5 years ago
Mass per unit area= M 1/2 ah
base= a
Moment of inerta= I = 1/2 x M/1/2ah x 1/2ax x xdx
x is the small strip along h from base
So, moment of inertia at base = 1/2 x M/h limit x².dx 
= M/2h (x³/ 3) 
= M/2h . h³/3 
= Mh²/6